OpenStudy (anonymous):

what would this be (sq rt 0f 21) * (sq rt 0f 3)

OpenStudy (anonymous):

Do like the last problem and multiply the two numbers under the root together and then find any perfect squares.

OpenStudy (anonymous):

how there are no perfect roots there is 7 and 3 put neight of those are perfect squaer

OpenStudy (anonymous):

There is 7*3=21 and 3 so your factors are 7,3,3.... because the problem is \[\sqrt{21*3}\]

OpenStudy (anonymous):

both problems have the same root so you can multiply the numbers under the root together.

OpenStudy (anonymous):

im still not getting this

OpenStudy (anonymous):

What do you get if you multiply 21*3? Then what are the factors of that? Then take out any factor pairs and what is left is what stays under the square root.

OpenStudy (anonymous):

63

OpenStudy (anonymous):

1and63 3and 21 9and 7 so i would use this one

OpenStudy (anonymous):

Right, because 9 is a perfect square and 7 doesn't have any perfect square factors.

OpenStudy (anonymous):

then what do i do with the other half of the problem

OpenStudy (anonymous):

umm what other half.... the problem was \[\sqrt{21}*\sqrt{3}\] correct?

OpenStudy (anonymous):

Read my second post.

OpenStudy (anonymous):

ya but what do i do with the (sq rt of 3)